Bill deposits
$700
into an account that pays simple interest at a rate of
3%
per year. How much interest will he be paid in the first
4
years?

Found 2 solutions by Alan3354, stanbon:
Answer by Alan3354(69443)

Bill deposits $700 into an account that pays simple interest at a rate of
3% per year. How much interest will he be paid in the first 4 years?
--------------
Find 3% of $700.
He gets that every year, so multiply it by 4.

Answer by stanbon(75887)

Bill deposits $700into an account that pays simple interest at a rate of 3%per year.
How much interest will he be paid in the first 4 years?
----
Interest = P*r*t = 700*0.03*4 = 700*0.12 = 12*7 = $84.00
---------
Cheers,
Stan H.
--------------